# EchoJEPA: A Latent Predictive Foundation Model for Echocardiography

### v1

@@ Abstract
Foundation models for echocardiography promise to reduce annotation burden and improve diagnostic consistency by learning generalizable representations from large unlabeled video archives. However, current approaches fail to disentangle anatomical signal from the stochastic speckle and acquisition artifacts that dominate ultrasound imagery. We present EchoJEPA, a foundation model for echocardiography trained on 18 million echocardiograms across 300K patients, the largest pretraining corpus for this modality to date. We also introduce a novel multi-view probing framework with factorized stream embeddings that standardizes evaluation under frozen backbones. Compared to prior methods, EchoJEPA reduces left ventricular ejection fraction estimation error by 19% and achieves 87.4% view classification accuracy. EchoJEPA exhibits strong sample efficiency, reaching 78.6% accuracy with only 1% of labeled data versus 42.1% for the best baseline trained on 100%. Under acoustic perturbations, EchoJEPA degrades by only 2.3% compared to 16.8% for the next best model, and transfers zero-shot to pediatric patients with 15% lower error than the next best model, outperforming all fine-tuned baselines. These results establish latent prediction as a superior paradigm for ultrasound foundation models.
@@ 1 Introduction
We introduce EchoJEPA, the first foundation-scale application of joint-embedding predictive architectures to echocardiography, pretrained on 18 million videos across 300K patients. EchoJEPA adapts V-JEPA2 with domain-appropriate temporal resolution and augmentation (Assran et al., 2025). To handle variable study composition, we develop a multi-view probing framework with factorized video stream embeddings that integrates information across views without view-specific components, offering a superior alternative to prior multi-view embedding scheme (Tohyama et al., 2025). Our results show that in medical domains where pixel fidelity is neither achievable nor desirable, objective-domain alignment determines representation quality more than architectural scale alone.
@@
  • Robustness benchmarks. Physics-informed perturbations using depth attenuation and acoustic shadow (Singla et al., 2022), revealing that EchoJEPA degrades 40% less than reconstruction-based baselines like VideoMAE (Tong et al., 2022).
  • Public release. We open-source EchoJEPA-L, a state-of-the-art echocardiography foundation model trained on MIMIC-IV-Echo (Gow et al., 2023), alongside our evaluation framework.
@@ 2.2 Foundation Models for Echocardiography
Public benchmarks catalyzed modern echocardiography learning. EchoNet-Dynamic (Ouyang et al., 2020) released 10,030 annotated apical-4-chamber videos and established supervised baselines for ejection fraction estimation, and EchoNet-Pediatric (Reddy et al., 2023) extended this to pediatric populations. Recent foundation models pursue broader task coverage through self-supervised pretraining. PanEcho (Holste et al., 2025) trains a ConvNeXt encoder with temporal transformer on over one million videos, achieving view-agnostic prediction through post-hoc averaging of clip-level outputs. EchoPrime (Vukadinovic et al., 2025) scales to 12 million video-report pairs with contrastive vision-language learning, using view classification and attention-based pooling for study-level predictions. Similarly, EchoCLIP (Christensen et al., 2023) aligns visual representations with text reports via contrastive learning on 1 million video-text pairs. EchoFM applies masked autoencoding to 290,000 videos to learn spatiotemporal representations (Kim et al., 2025).
@@ 2.3 Robustness Evaluation in Medical Imaging
Standard evaluation protocols emphasize i.i.d. test performance on held-out splits from the training distribution, underestimating failure modes in clinical deployment where distribution shift is the norm (Quionero-Candela et al., 2009). Echocardiography exhibits acquisition variability since image quality depends on patient body habitus, probe positioning, operator expertise, and equipment vendor (Picard et al., 2011). Patients most likely to benefit from automated analysis, such as those with obesity or limited acoustic windows, are precisely those whose images deviate most from training distributions. Prior robustness evaluations focus on ImageNet-C corruptions (Hendrycks and Dietterich, 2019) or adversarial perturbations (Finlayson et al., 2019), neither capturing ultrasound-specific degradation.
Figure 1: EchoJEPA architecture. Echocardiograms from multiple views are partitioned into spatio-temporal tubelets and split into masked and unmasked sets. The encoder $E_{\theta}$ processes masked video frames, and the predictor $P_{\phi}$ infers embeddings for masked regions conditioned on learnable mask tokens. The EMA encoder $E_{\bar{\theta}}$ processes unmasked frames to provide prediction targets. The $L_{1}$ loss is computed between predicted and target embeddings, with no gradients flowing into the EMA encoder.
@@ 4.1 Experimental Setup
  • EchoJEPA-L: ViT-Large (307M params) trained with latent prediction on 525K videos from MIMIC-IV-Echo (Gow et al., 2023) (public).
@@ 4.3 Sample Efficiency
We evaluate sample efficiency by training probes on 1%, 10%, and 100% of 22,000 labeled views. (Table 4). Remarkably, both EchoJEPA models trained on just 1% of labeled data outperform all reconstruction baselines trained on 100% of the data. EchoJEPA-L, despite being trained on public data, achieves 57.6% accuracy with 1% labels, surpassing the best performing baseline (EchoPrime at 42.1%) by a wide margin. EchoJEPA-G extends this lead further, reaching 78.6% accuracy with 1% labels, nearly double the performance of fully-supervised baselines.
@@ 4.4 Robustness to Acoustic Degradation
Table 5: Robustness to acoustic degradation (LVEF MAE on Stanford, lower is better). EchoJEPA degrades more gracefully than reconstruction-based models under depth attenuation and Gaussian shadowing. Avg. Deg. reports relative increase from clean performance.
@@ 4.5 Multi-View Physiological Estimation
Table 6 reports RVSP estimation, which requires integrating apical and subcostal views. EchoJEPA-G achieves 4.54 MAE on Toronto, an 18% improvement over PanEcho.
@@ 5 Discussion
EchoJEPA outperforming compute-matched reconstruction baselines challenges the assumption that methods from natural video transfer directly to medical imaging. In natural video, texture correlates with semantics; in ultrasound, texture is largely interference noise. Pixel reconstruction forces the model to memorize this noise, while latent prediction captures only spatiotemporally coherent structures, yielding representations that are over 85% more robust to acoustic perturbations.
@@
The proposed multi-view probing framework resolves a long-standing evaluation challenge. Prior works (Holste et al., 2025; Vukadinovic et al., 2025) relied on disparate protocols, making rigorous comparison impossible. Our framework handles variable study composition without view-specific encoders, establishing a baseline for future research.
@@
While we release EchoJEPA-L trained on public MIMIC-IV-Echo data, our strongest model relies on proprietary data that cannot be released. Additionally, our robustness evaluation uses synthetic perturbations, and prospective validation on difficult-to-image patients remains the gold standard.
@@
EchoJEPA opens directions including fine-grained tasks such as valve segmentation, prospective validation on patients with poor acoustic windows, and extension to other noisy modalities such as fetal and lung ultrasound.
@@ 6 Conclusion
We introduce EchoJEPA, a foundation model for echocardiography demonstrating that latent prediction outperforms pixel reconstruction and alternative pretraining objectives. Trained on 18 million videos across 300K patients, EchoJEPA achieves state-of-the-art performance on LVEF estimation, RVSP prediction, and view classification, reducing LVEF error by 27% over compute-matched reconstruction baselines. EchoJEPA reaches 78.6% view accuracy with 1% of labels versus 42.1% for the best baseline at 100%, degrades by only 2.3% under acoustic perturbations compared to 17% for alternatives, and transfers zero-shot to pediatric patients better than fine-tuned baselines. We release EchoJEPA-L trained on public MIMIC-IV-Echo data alongside our multi-view probing framework.
@@
This work aims to improve automated echocardiography analysis, which could expand access to expert-level cardiac assessment in resource-limited settings. Sample-efficient and robust models may particularly benefit patients who are currently underserved—those with obesity, lung disease, or limited access to trained cardiologists. However, automated cardiac assessment deployed without adequate validation could lead to diagnostic errors, and models trained on data from high-resource healthcare systems may underperform in other clinical contexts. We release EchoJEPA-L on public data to enable independent evaluation, explicitly characterize limitations, and emphasize that our models are research artifacts requiring clinical validation before deployment.
@@ Acknowledgements
We are grateful to members of the Meta AI team, particularly Quentin Garrido and Koustuv Sinha, for their invaluable guidance in understanding JEPA training dynamics and adapting the architecture for the echocardiography domain. We thank Augustin Toma and Jun Ma for insightful discussions on adapting the model to multi-view clinical tasks through early fusion and cross-attention probing mechanisms. We gratefully acknowledge Amazon Web Services for providing the computational infrastructure and technical support necessary for training and evaluating models at this scale; we especially thank Joshua Thomas and Elizabeth Keller for their partnership in advancing AI for healthcare. Finally, we thank the University Health Network for providing the institutional resources that made this research possible.
Appendix A Compute-Matched Pretraining Protocol
@@ A.1 Shared Configuration
We sample 16 frames per clip at 8 FPS ($\approx$2.0 seconds temporal coverage), tokenized with patch size 16 and tubelet size 2. Spatial augmentation uses random resized crops with scale $[0.5,1.0]$ and aspect ratio $[0.9,1.1]$, narrower than the V-JEPA2 defaults to preserve clinically meaningful chamber geometry (see Section 3.2). Inputs are normalized with ImageNet statistics.
@@
  • Total optimizer updates: 72,000 (pretraining) + 12,000 (cooldown) = 84,000
@@
A.2 V-JEPA Configuration (EchoJEPA-L)
@@
Architecture ViT-Large (307M params)
@@
A.3 VideoMAE Configuration (EchoMAE-L)
@@
Architecture ViT-Large (307M params)
@@
Learning rate $8.79\times 10^{-7}$ (scaled by batch size)
@@ A.5 Summary of Controlled Variables
Appendix B Ultrasound-Specific Data Augmentation
@@ B.1 Overview of USAugment
These transforms integrate seamlessly with the Albumentations framework (albumentations2020) and require a binary scan mask $M\in\{0,1\}^{H\times W}$ that identifies the active ultrasound scan region within the image frame. This mask ensures augmentations are applied only to the diagnostic region, preserving any surrounding interface elements or annotations.
B.2 Scan Mask Generation
@@
$M_{\text{sector}}(x,y)=\begin{cases}1&\text{if }\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}\leq r_{\max}\text{ and }\left|\arctan\left(\frac{x-x_{0}}{y-y_{0}}\right)\right|\leq\theta\\ 0&\text{otherwise}\end{cases}$ (7)
@@
B.3 Depth Attenuation
B.3.1 Physical Motivation
@@ B.3.4 Experimental Configurations
B.4 Gaussian Shadow
B.4.1 Physical Motivation
@@
The Gaussian shadow transform, originally described by Smistad et al. (smistad2018) for nerve identification in ultrasound, provides a computationally efficient approximation of these acoustic shadows using 2D Gaussian functions.
B.4.2 Implementation
@@
B.4.3 Temporal Consistency for Video
@@ B.4.5 Experimental Configurations
B.5 Haze Artifact
B.5.1 Physical Motivation
@@ B.5.3 Parameters
B.6 Speckle Reduction
B.6.1 Physical Motivation
@@ B.6.3 Parameters
B.7 Composing Multiple Augmentations
@@
B.8 Augmentation Strategy for Video Models
@@ B.9 Software and Reproducibility
Appendix C Ablation Studies

### v2

@@ Abstract
Foundation models for echocardiography often struggle to disentangle anatomical signal from the stochastic speckle and acquisition artifacts inherent to ultrasound. We present EchoJEPA, a foundation model trained on 18 million echocardiograms across 300K patients, representing the largest pretraining corpus for this modality to date. By leveraging a latent predictive objective, EchoJEPA learns robust anatomical representations that ignore speckle noise. We validate this using a novel multi-view probing framework with frozen backbones, where EchoJEPA outperforms state-of-the-art baselines by approximately 20% in left ventricular ejection fraction (LVEF) estimation and 17% in right ventricular systolic pressure (RVSP) estimation. The model also exhibits remarkable sample efficiency, reaching 79% view classification accuracy with only 1% of labeled data versus 42% for the best baseline trained on 100%. Crucially, EchoJEPA demonstrates superior generalization, degrading by only 2% under physics-informed acoustic perturbations compared to 17% for competitors. Most remarkably, its zero-shot performance on pediatric patients surpasses fully fine-tuned baselines, establishing latent prediction as a superior paradigm for robust, generalizable medical AI.
@@ 1 Introduction
We introduce EchoJEPA, the first foundation-scale application of joint-embedding predictive architectures to echocardiography, pretrained on 18 million videos across 300K patients. EchoJEPA adapts V-JEPA2 with domain-appropriate temporal resolution and augmentation (Assran et al., 2025). To handle variable study composition, we develop a multi-view probing framework with factorized video stream embeddings that integrates information across views without view-specific components, offering a superior alternative to prior multi-view embedding scheme (Tohyama et al., 2025).
@@
  • Robustness benchmarks. Physics-informed perturbations using depth attenuation and acoustic shadow (Singla et al., 2022), revealing that EchoJEPA degrades 86% less than the next-best baseline under acoustic perturbations.
  • Public release. We open-source EchoJEPA-L, a state-of-the-art echocardiography foundation model trained on MIMIC-IV-Echo (Gow et al., 2023), alongside our evaluation framework at https://github.com/bowang-lab/EchoJEPA.
@@ 2.2 Foundation Models for Echocardiography
Building on foundation model success in healthcare, echocardiography has followed suit (Ma et al., 2024, 2025; Fallahpour et al., 2024). EchoNet-Dynamic (Ouyang et al., 2020) released 10,030 annotated apical-4-chamber videos and established supervised baselines for ejection fraction estimation, and EchoNet-Pediatric (Reddy et al., 2023) extended this to pediatric populations. Recent foundation models pursue broader task coverage through self-supervised pretraining. PanEcho (Holste et al., 2025) trains a ConvNeXt encoder with temporal transformer on over one million videos, achieving view-agnostic prediction through post-hoc averaging of clip-level outputs. EchoPrime (Vukadinovic et al., 2025) scales to 12 million video-report pairs with contrastive vision-language learning, using view classification and attention-based pooling for study-level predictions. Similarly, EchoCLIP (Christensen et al., 2023) aligns visual representations with text reports via contrastive learning on 1 million video-text pairs. EchoFM applies masked autoencoding to 290,000 videos to learn spatiotemporal representations (Kim et al., 2025).
@@ 2.3 Robustness Evaluation in Medical Imaging
Standard evaluation protocols emphasize i.i.d. test performance, underestimating failure modes in clinical deployment where distribution shift is ubiquitous (Quionero-Candela et al., 2009). Echocardiography exhibits acquisition variability since image quality depends on patient body habitus, probe positioning, operator expertise, and equipment vendor (Picard et al., 2011). Patients most likely to benefit from automated analysis, such as those with obesity or limited acoustic windows, are those whose images deviate most from training distributions, yet prior robustness evaluations focus on ImageNet-C corruptions (Hendrycks and Dietterich, 2019) or adversarial perturbations (Finlayson et al., 2019), neither capturing ultrasound-specific degradation.
Figure 1: EchoJEPA architecture. Echocardiograms from multiple views are partitioned into spatio-temporal tubelets and split into masked and unmasked sets. The encoder $E_{\theta}$ processes visible (unmasked) video frames, and the predictor $P_{\phi}$ infers embeddings for masked regions conditioned on learnable mask tokens. The EMA encoder $E_{\bar{\theta}}$ processes unmasked frames to provide prediction targets. The $L_{1}$ loss is computed between predicted and target embeddings, with no gradients flowing into the EMA encoder.
@@ 4.1 Experimental Setup
  • EchoJEPA-L: ViT-Large (300M params) trained with latent prediction on 525K videos from MIMIC-IV-Echo (Gow et al., 2023) (public).
@@ 4.3 Sample Efficiency
Table 4 demonstrates that EchoJEPA models trained on just 1% of labeled data outperform baselines trained on 100%. Even the publicly trained EchoJEPA-L achieves 57.6% accuracy with 1% labels (vs. 42.1% for EchoPrime), while EchoJEPA-G reaches 78.6%, nearly double the fully-supervised baseline. This efficiency implies that latent prediction yields dense representations capable of defining the view manifold with minimal supervision, as evidenced by the distinct anatomical clustering in Figure 5.
@@ 4.4 Robustness to Acoustic Degradation
Table 5: Robustness to acoustic degradation (LVEF MAE on Stanford, lower is better). EchoJEPA degrades more gracefully than alternative foundation models under depth attenuation and Gaussian shadowing. Avg. Deg. reports relative increase from clean performance.
@@ 4.5 Multi-View Physiological Estimation
Table 6 reports RVSP estimation, which requires integrating apical and subcostal views. EchoJEPA-G achieves 4.54 MAE on Toronto, a 17% improvement over PanEcho.
@@ 5 Discussion
EchoJEPA outperforming compute-matched reconstruction baselines challenges the assumption that methods from natural video transfer directly to medical imaging. In natural video, texture correlates with semantics; in ultrasound, texture is largely interference noise. Pixel reconstruction forces the model to memorize this noise, while latent prediction captures only spatiotemporally coherent structures, yielding representations up to over 85% more robust to acoustic perturbations.
@@
Our multi-view probing framework resolves a crucial evaluation challenge. Prior works (Holste et al., 2025; Vukadinovic et al., 2025) relied on disparate protocols, making rigorous comparison impossible. Our framework handles variable study composition without view-specific encoders, a baseline for future research.
@@
Primary limitations include the reliance on proprietary data for our strongest model (though EchoJEPA-L is public), the use of synthetic rather than prospective clinical perturbations, and potential privacy risks regarding memorization that warrant further study (Tonekaboni et al., 2025).
@@
EchoJEPA opens directions including fine-grained tasks such as valve segmentation, prospective validation on patients with poor acoustic windows, integration with interpretable reasoning frameworks for clinical decision support (Fallahpour et al., 2025), and extension to other noisy modalities such as fetal and lung ultrasound.
@@ 6 Conclusion
We introduce EchoJEPA, a foundation model for echocardiography demonstrating that latent prediction outperforms pixel reconstruction and alternative pretraining objectives. Trained on 18 million videos across 300K patients, EchoJEPA achieves state-of-the-art performance on LVEF estimation, RVSP prediction, and view classification, reducing LVEF error by 27% over compute-matched reconstruction baselines. EchoJEPA reaches 78.6% view accuracy with 1% of labels versus 42.1% for the best baseline at 100%, degrades by only 2.3% under acoustic perturbations compared to 16.8% for the next-best foundation model, and transfers zero-shot to pediatric patients better than fine-tuned baselines. We release EchoJEPA-L trained on public MIMIC-IV-Echo data alongside our multi-view probing framework.
@@
This work aims to improve automated echocardiography analysis, which could expand access to expert-level cardiac assessment in resource-limited settings. Sample-efficient and robust models may particularly benefit patients who are currently underserved—those with obesity, lung disease, or limited access to trained cardiologists. However, automated cardiac assessment deployed without adequate validation could lead to diagnostic errors, and models trained on data from high-resource healthcare systems may underperform in other clinical contexts. We release EchoJEPA-L trained on public data to enable independent evaluation, explicitly characterize limitations, and emphasize that our models are research artifacts requiring clinical validation before deployment.
@@ Acknowledgements
We extend our gratitude to Amazon Web Services for providing the computational infrastructure essential to this work, and specifically to Joshua Thomas and Elizabeth Keller for their partnership in advancing AI for healthcare. We are particularly indebted to Quentin Garrido and Koustuv Sinha of the Meta AI team for their invaluable guidance on JEPA training dynamics and architectural adaptation. We also thank Augustin Toma and Jun Ma for insightful discussions regarding multi-view early fusion and cross-attention probing, as well as Zhibin Lu for his expertise in distributed systems and data pipelines. Finally, we acknowledge the University Health Network for the institutional resources that made this research possible.
Appendix A Latent Space Analysis
To assess the semantic quality of learned representations, we visualize the frozen embedding spaces of all evaluated models using Uniform Manifold Approximation and Projection (UMAP).
As shown in Figure 5, EchoJEPA-G demonstrates strong semantic organization, forming distinct, well-separated clusters for different anatomical views (e.g., PLAX, A4C). Notably, the model clearly segregates Transesophageal (TEE) echocardiograms into a discrete cluster separate from standard Transthoracic (TTE) views. This indicates that latent prediction effectively disentangles acquisition modalities without explicit supervision.
In contrast, baselines such as EchoMAE-L (reconstruction), EchoPrime (contrastive), and PanEcho (supervised) exhibit diffuse embedding spaces where TTE and TEE views are largely intermixed. This suggests that alternative objectives fail to capture fundamental anatomical distinctions as effectively as latent prediction. The visual clustering quality observed here strongly correlates with the probe accuracy reported for each model.
Figure 5: UMAP visualization of frozen video representations colored by echocardiographic view. Baselines (left) exhibit diffuse distributions with significant class overlap, correlating with lower probe accuracy. EchoJEPA models (right) form distinct anatomical clusters, including a clear separation of Transesophageal (TEE) views.
Appendix B Compute-Matched Pretraining Protocol
@@ B.1 Shared Configuration
We sample 16 frames per clip at 8 FPS (distinct from the 24 FPS used for the proprietary EchoJEPA-G to match the standard VideoMAE protocol), tokenized with patch size 16 and tubelet size 2. Spatial augmentation uses random resized crops with scale $[0.5,1.0]$ and aspect ratio $[0.9,1.1]$, narrower than the V-JEPA2 defaults to preserve clinically meaningful chamber geometry (see Section 3.2). Inputs are normalized with ImageNet statistics, tokenized with patch size 16 and tubelet size 2.
@@
  • Total optimizer updates: 60,000 (pretraining) + 24,000 (cooldown) = 84,000
@@
B.2 V-JEPA Configuration (EchoJEPA-L)
@@
Architecture ViT-Large (300M params)
@@
B.3 VideoMAE Configuration (EchoMAE-L)
@@
Architecture ViT-Large (300M params)
@@
Effective Learning Rate $3.52\times 10^{-6}$
@@ B.4 Initialization
Appendix C Ultrasound-Specific Data Augmentation
@@ C.1 Overview of USAugment
These transforms integrate seamlessly with the Albumentations framework (Buslaev et al., 2020) and require a binary scan mask $M\in\{0,1\}^{H\times W}$ that identifies the active ultrasound scan region within the image frame. This mask ensures augmentations are applied only to the diagnostic region, preserving any surrounding interface elements or annotations.
C.2 Scan Mask Generation
@@
$M_{\text{sector}}(x,y)=\begin{cases}1&\text{if }\sqrt{(x-x_{0})^{2}+(y-y_{0})^{2}}\leq r_{\max}\quad\text{and}\quad\left|\operatorname{atan2}(x-x_{0},y-y_{0})\right|\leq\theta\\ 0&\text{otherwise}\end{cases}$ (7)
@@
C.3 Depth Attenuation
C.3.1 Physical Motivation
@@ C.3.4 Experimental Configurations
C.4 Gaussian Shadow
C.4.1 Physical Motivation
@@
The Gaussian shadow transform, originally described by Smistad et al. (Smistad et al., 2018) for nerve identification in ultrasound, provides a computationally efficient approximation of these acoustic shadows using 2D Gaussian functions.
C.4.2 Implementation
@@
C.4.3 Temporal Consistency for Video
@@ C.4.5 Experimental Configurations
C.5 Haze Artifact
C.5.1 Physical Motivation
@@ C.5.3 Parameters
C.6 Speckle Reduction
C.6.1 Physical Motivation
@@ C.6.3 Parameters
C.7 Composing Multiple Augmentations
@@
C.8 Augmentation Strategy for Video Models
@@ C.9 Software and Reproducibility
Appendix D Ablation Studies